# Intrinsic Wrapped Gaussian Process Regression Modeling for Manifold-valued Response Variable

### v1

@@ 2.2 Intrinsic covariance function of tangent space
Sphere manifold $\mathcal{S}^{D}$. Suppose that $\mu(t)$ is a geodesic curve. The parallel transformation $\Gamma_{\mu(t)\rightarrow\mu(s)}(\boldsymbol{\gamma}_{t})$ is $\mathcal{A}_{\mu(t)\rightarrow\mu(s)}\boldsymbol{\gamma}_{t}$, where $\mathcal{A}_{\mu(t)\rightarrow\mu(s)}$ is a transformation matrix and $\boldsymbol{\gamma}_{t}\in T_{\mu(t)}\mathcal{S}^{D}$. Then, by Definition 2.1 and Proposition 2.2, we have
$\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}(\cdot)=\mathcal{A}_{\mu(t_{1})\rightarrow\mu(t_{2})}(\cdot)\mathcal{A}^{\top}_{\mu(t_{1})\rightarrow\mu(t_{2})},$ (4)
$\boldsymbol{\Phi}_{\mu(s)\rightarrow\mu(t)}(\cdot)=\mathcal{A}_{\mu(s)\rightarrow\mu(t)}(\cdot)\mathcal{A}^{\top}_{\mu(s)\rightarrow\mu(t)}.$
@@
SPD manifold $\mathcal{S}_{++}^{D}$ with affine-invariant metric. Suppose that $\mu(t)$ is a geodesic curve. The parallel transformation $\Gamma_{\mu(t)\rightarrow\mu(s)}(\boldsymbol{\gamma}_{t})$ is $\mathcal{R}_{\mu(t)\rightarrow\mu(s)}\boldsymbol{\gamma}_{t}\mathcal{R}_{\mu(t)\rightarrow\mu(s)}^{\top}$, where $\boldsymbol{\gamma}_{t}\in T_{\mu(t)}\mathcal{S}_{++}^{D}$ is asymmetric matrix and $\mathcal{R}_{\mu(t)\rightarrow\mu(s)}=\mu(s)^{1/2}\mu(t)^{-1/2}$. Then, by Definition 2.1 and Proposition 2.2, we have
$\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}(\cdot)=(\mathcal{R}_{\mu(t_{1})\rightarrow\mu(t_{2})}\otimes\mathcal{R}_{\mu(t_{1})\rightarrow\mu(t_{2})})(\cdot)(\mathcal{R}^{\top}_{\mu(t_{1})\rightarrow\mu(t_{2})}\otimes\mathcal{R}^{\top}_{\mu(t_{1})\rightarrow\mu(t_{2})}),$ (5)
$\boldsymbol{\Phi}_{\mu(s)\rightarrow\mu(t)}(\cdot)=(\mathcal{R}_{\mu(s)\rightarrow\mu(t)}\otimes\mathcal{R}_{\mu(s)\rightarrow\mu(t)})(\cdot)(\mathcal{R}^{\top}_{\mu(s)\rightarrow\mu(t)}\otimes\mathcal{R}^{\top}_{\mu(s)\rightarrow\mu(t)}),$
where $\otimes$ denotes tensor product.
The proof of examples 2.4-2.5 are given in Appendix C. For other manifolds, such as Hyperbolic manifold and Grassmannian manifold [28], the parallel transformation $\boldsymbol{\Phi}$ can be derived similarly.
@@
where $\mu(\cdot)$ is the basis point function (BPF) on a Riemannian manifold, and $k_{t}(\cdot,\cdot)$ defined in (2) or (3) is an intrinsic covariance function on a tangent space at $\mu(t)$. In this paper, we call (6) an intrinsic wrapped Gaussian distribution with the basepoint vector $\boldsymbol{\mu}_{n}$ and the covariance matrix $K$.
2.3 Intrinsic wrapped Gaussian process regression
Using the intrinsic wrapped Gaussian process, we define an intrinsic wrapped Gaussian process regression (iWGPR) model for the manifold-valued response $y\in\mathcal{M}$ and the vector-valued predictor $\boldsymbol{x}\in\Omega$ via an unknown regression function $f:\Omega\rightarrow\mathcal{M}$. Let $\Omega=\mathbb{R}^{Q}$ firstly. Later, we will extend $\Omega$ to other spaces. Given the training data $\mathcal{D}_{\mathcal{M}}:=\left\{\left(\boldsymbol{x}_{t_{i}},y_{t_{i}}\right)\mid\boldsymbol{x}_{t_{i}}\in\right.\left.\mathbb{R}^{Q},y_{t_{i}}\in\mathcal{M},i=1,\cdots,n,t_{i}\in\mathbb{T}\right\}$, where $\mathbb{T}$ is a compact set, the main idea of the model is to assume an intrinsic wrapped Gaussian process as the prior of $f(\cdot)$. We consider ${T}_{\mu(t^{*})}\mathcal{M}$ as the common tangent space. We will show later that the estimation results are invariant to the choice of this common tangent space. The joint distribution between the training outputs $\boldsymbol{y}=\left(y_{t_{1}},\cdots,y_{t_{n}}\right)$ and the test outputs $y_{t^{*}}$ at $\boldsymbol{x}_{t^{*}}$ is given by
@@ 2.3 Intrinsic wrapped Gaussian process regression
where $\mu^{*}=\mu\left(t^{*}\right)$, $\boldsymbol{\mu}=\left(\mu\left(t_{1}\right),\ldots,\mu\left(t_{{n}}\right)\right)$, $\boldsymbol{X}=\left(\boldsymbol{x}_{t_{1}},\cdots,\boldsymbol{x}_{t_{n}}\right)$, $K_{t^{*}}\left(\boldsymbol{X},\boldsymbol{x}^{*}\right)=(k_{t^{*}}\left(\boldsymbol{x}_{t_{1}},\boldsymbol{x}^{*}\right),k_{t^{*}}\left(\boldsymbol{x}_{t_{2}},\boldsymbol{x}^{*}\right),\cdots,k_{t^{*}}\left(\boldsymbol{x}_{t_{n}},\boldsymbol{x}^{*}\right))$ and $K_{t^{*}}(\boldsymbol{X},\boldsymbol{X})=\{k_{t^{*}}(\boldsymbol{x}_{t_{i}},\boldsymbol{x}_{t_{j}})\}_{1\leq i,j\leq n}$ and $k_{t^{*}}(\cdot,\cdot)$ is defined in (7), but with the common tangent space at $\mu(t^{*})$.
@@
In practice, we might need to work with a specific orthonormal bases for the tangent space. Therefore, we pick an orthogonal coordinate frame $\mathbf{E}$ for the common tangent space ${T}_{\mu(t^{*})}\mathcal{M}$, and for any $t\in\mathbb{T}$, $\mathbf{E}_{t}=\Gamma_{\mu(t^{*})\rightarrow\mu(t)}(\mathbf{E})$ is an orthogonal coordinate frame for the tangent space ${T}_{\mu(t)}\mathcal{M}$. Next, we will also show that the model estimation results are invariant to the choice of the orthogonal coordinate frame. Let $k_{\mathbf{E},t^{*}}$ be the representation of the intrinsic tangent space covariance function $k_{t^{*}}$ in frame $\mathbf{E}$, and $\text{cov}_{\mathbf{E}}$ be the representation of the covariance function in frame $\mathbf{E}$. Thence, we have $k_{t^{*}}=\mathbf{E}^{\top}k_{\mathbf{E},t^{*}}\mathbf{E}$. And for any $t_{1},t_{2}\in\mathbf{T}$, we have
$\text{cov}_{\mathbf{E}}(\Gamma_{\mu(t_{1})\rightarrow\mu(t^{*})}(\boldsymbol{\gamma}_{t_{1}}),\Gamma_{\mu(t_{2})\rightarrow\mu(t^{*})}(\boldsymbol{\gamma}_{t_{2}})):=k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}}).$ (9)
Note that $k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}})$ is the matrix-valued function. To obtain a valid $k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}})$, a commonly used method is the intrinsic coregionalization model. This leads to
$k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}}):=k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}},\boldsymbol{\theta})\otimes\mathbf{B}\quad\text{with}\quad\mathbf{B}\in\mathbf{R}^{D\times D},$ (10)
where $k_{\mathbb{R}^{Q}}$ is a scalar kernel function, $\boldsymbol{\theta}$ is its kernel parameters, and $\mathbf{B}$ is a $D\times D$ coregionalization matrix that is positive semi-definite. Let matrix $K_{t^{*}}=(k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t_{i}},\boldsymbol{x}_{t_{j}},\boldsymbol{\theta}))_{n\times n}$, and $\boldsymbol{k}_{t^{*}}=(k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t^{*}},\boldsymbol{x}_{t_{1}},\boldsymbol{\theta}),k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t^{*}},\boldsymbol{x}_{t_{2}},\boldsymbol{\theta}),\cdots,k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t^{*}},\boldsymbol{x}_{t_{n}},\boldsymbol{\theta}))^{\top}$. Then predictive distribution of the test data $\boldsymbol{x}_{t^{*}}$ in tangent space of $\mu(t^{*})$ conditional on the training data $\mathcal{D}_{\mathcal{M}}$ is obtained as
@@
It follows that $\operatorname{Exp}(\mu(t^{*}),{\boldsymbol{\beta}}_{t^{*}}^{*})$ can be used as a point prediction of $y_{t^{*}}$. Although the predictive distribution $\mathcal{N}_{t^{*}}$ is not necessarily WGD, as $\boldsymbol{\beta}_{t^{*}}^{*}$ might be non-zero. As in [24], the distribution can be approximated by using a WGD with the basepoint $\operatorname{Exp}_{{\mu}(t^{*})}\left(\boldsymbol{\beta}_{t^{*}}^{*}\right)$ and parallel transporting the tangent space covariance $\boldsymbol{\Sigma}_{t^{*}}^{*}$ to this basepoint along the corresponding geodesic. In addition, these hyperparameters $\boldsymbol{\theta}$ and $\mathbf{B}$ can be obtained by maximizing the log marginal likelihood. As there is no analytic solution, the recursive algorithm developed by [31, 32] can be used. A detailed computation procedure for iWGPR is proved in Algorithm 1.
Algorithm 1 iWGPR
@@
Next, we need to address the problem of how to choose the common tangent space. The process so far has been carried out in the tangent space of $\mu(t^{*})$. What would happen if we were to perform the analysis in the tangent space of other points along the BPF ? Theorem 2.6 below provides an answer: the posterior distribution $\mathcal{N}_{t}$ under the tangent space of other points $\mu(t)$ on the BPF, after applying a parallel transformation $\boldsymbol{\Phi}_{\mu(t)\rightarrow\mu(t^{*})}$ along the BPF, remains consistent with the posterior distribution $\mathcal{N}_{t^{*}}$ in the tangent space of $\mu(t^{*})$. The proof is given in Appendix D.
@@
Assuming that the BPF $\mu(t)$ is a smooth curve. For any $t\in\mathbf{T}$, let $\mathbf{A}$ be a frame for the tangent space of $\mu(t)$ and $\Gamma_{\mu(t)\rightarrow\mu(t^{*})}(\mathbf{A})=\mathbf{E}$. Then, we have
${\boldsymbol{\beta}}_{t^{*}}^{*}=\Gamma_{\mu(t)\rightarrow\mu(t^{*})}({\boldsymbol{\beta}}_{t}^{*});\quad\boldsymbol{\Sigma}_{t^{*}}^{*}=\boldsymbol{\Phi}_{\mu(t)\rightarrow\mu(t^{*})}(\boldsymbol{\Sigma}_{t}^{*})$ (12)
where ${\boldsymbol{\beta}}_{t}^{*}$ and $\boldsymbol{\Sigma}_{t}^{*}$ represent the mean and variance of the predicted distribution $\mathcal{N}_{t}$ in the tangent space of $\mu(t)$.
Furthermore, the following Theorem 2.7 demonstrates that the predicted distribution is independent of the chosen frame. The proof is given in Appendix E.
@@
For any $t\in\mathbf{T}$, suppose $\mathbf{W}$ is another frame on the tangent spaces of $\mu(t)$. Change from $\mathbf{E}$ to $\mathbf{W}$ can be characterized by a unitary matrix $\mathbf{O}$, and $\mathbf{W}=\mathbf{O}^{\top}\mathbf{E}$. Then, we have
${\boldsymbol{\beta}}_{\mathbf{E},t}^{*}=\mathbf{O}{\boldsymbol{\beta}}_{\mathbf{W},t}^{*};\quad\overline{\boldsymbol{\Sigma}}^{*}_{\mathbf{E},t}=\mathbf{O}\overline{\boldsymbol{\Sigma}}^{*}_{\mathbf{W},t}\mathbf{O}^{\top}.$ (13)
This is, the mean and variance of the predicted distribution is independent of the frame.
Figure 2: Logarithm map of the manifold-valued data and schematic diagram of equivalent version of iWGPR, where the blue solid line represents BPF.
The proposed iWGPR utilities the curvature information of the manifold by implementing parallel transport from different tangent spaces along the BPF $\mu(t)$ to a common tangent space, affectively merging the intrinsic geometric structure of the manifold. We have shown that the estimation for the iWGPR model is unaffected by the selection of coordinate frames or common tangent space. In addition, we find an equivalent version of the proposed iWGPR: parallel transporting the coordinate frame along the BPF results in a series of coordinate frame across different tangent spaces, leading to a representation of random elements in these frames within a Euclidean space $\mathbb{R}^{D}$. We also show that these representations retain consistency despite changes in the tangent spaces. Figure 2 gives a geometric illustration, where the black arrows represent coordinate frames obtained via parallel transport operator on different tangent spaces, and the orange solid lines stand for the representation of random elements after a parallel transport from the common tangent space to different tangent spaces. This results in the purple dashed lines connecting $y_{i}$, which is the curve predicted from the intrinsic wrapped Gaussian process regression model. The relative position between the orange solid line and coordinate axes in black remains unchanged, indicating the choice of different common tangent space does not affect the model estimation. After representing the random elements of different tangent spaces using the aforementioned coordinate frame, they reside within the same Euclidean space and can be modelled by a vector-valued Gaussian process regression model.
We assumed Euclidean predictor so far. The proposed method however can be extended to a broader framework that can model response variables residing on manifolds with mixed Euclidean or non-Euclidean predictors.
@@
By incorporating kernel functions specifically designed for non-Euclidean predictors within the iWGPR model, our proposed method still work. Let the predictor $\boldsymbol{x}\in\Omega$, where $\Omega$ is a general space such as metric space. As long as there is kernel function $k_{\Omega}$ with symmetric and positive definite properties, the computations of estimation and prediction are carried out similar with those in Subsection 2.3 while $k_{\mathbb{R}^{Q}}$ is replaced with $k_{\Omega}$. For example, when $\Omega$ is a Riemannian manifold, one might select either the heat kernel [33, 34] or the Matérn kernel [12]. When $\Omega$ is a general metric space, options include graph kernel [35, 36] and persistence weighted Gaussian kernel [37].
2.4 iWGPR with observation noise
We didn’t consider noise in the model discussed in Section 2.3. In practical applications, data typically include random fluctuations and observational errors. By incorporating noise into the iWGPR framework, the model can more accurately distinguish between signal and noise and then capture the underlying patterns. Furthermore, the inclusion of noise guarantees the positive definiteness of the covariance matrix, ensuring numerical stability. The iWGPR model with error terms can be defined as

### v2

@@ 2.2 Intrinsic covariance function of tangent space
Sphere manifold $\mathcal{S}^{D}$. Suppose that $\mu(t)$ is a geodesic curve. The parallel transformation $\Gamma_{\mu(t)\rightarrow\mu(s)}(\boldsymbol{\gamma}_{t})$ is $\mathcal{A}_{\mu(t)\rightarrow\mu(s)}\boldsymbol{\gamma}_{t}$, where $\mathcal{A}_{\mu(t)\rightarrow\mu(s)}$ is a transformation matrix and $\boldsymbol{\gamma}_{t}\in T_{\mu(t)}\mathcal{S}^{D}$. Then, by Definition 2.1 and Proposition 2.2, we have $\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}(\cdot)=\mathcal{A}_{\mu(t_{1})\rightarrow\mu(t_{2})}(\cdot)\mathcal{A}^{\top}_{\mu(t_{1})\rightarrow\mu(t_{2})}$ and $\boldsymbol{\Phi}_{\mu(s)\rightarrow\mu(t)}(\cdot)=\mathcal{A}_{\mu(s)\rightarrow\mu(t)}(\cdot)\mathcal{A}^{\top}_{\mu(s)\rightarrow\mu(t)}$.
@@
SPD manifold $\mathcal{S}_{++}^{D}$ with affine-invariant metric. Suppose that $\mu(t)$ is a geodesic curve. The parallel transformation $\Gamma_{\mu(t)\rightarrow\mu(s)}(\boldsymbol{\gamma}_{t})$ is $\mathcal{R}_{\mu(t)\rightarrow\mu(s)}\boldsymbol{\gamma}_{t}\mathcal{R}_{\mu(t)\rightarrow\mu(s)}^{\top}$, where $\boldsymbol{\gamma}_{t}\in T_{\mu(t)}\mathcal{S}_{++}^{D}$ is an asymmetric matrix, and $\mathcal{R}_{\mu(t)\rightarrow\mu(s)}=\mu(s)^{1/2}\mu(t)^{-1/2}$. Then, by Definition 2.1 and Proposition 2.2, we have $\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}(\cdot)=(\mathcal{R}_{\mu(t_{1})\rightarrow\mu(t_{2})}\otimes\mathcal{R}_{\mu(t_{1})\rightarrow\mu(t_{2})})(\cdot)(\mathcal{R}^{\top}_{\mu(t_{1})\rightarrow\mu(t_{2})}\otimes\mathcal{R}^{\top}_{\mu(t_{1})\rightarrow\mu(t_{2})})$ and $\boldsymbol{\Phi}_{\mu(s)\rightarrow\mu(t)}(\cdot)=(\mathcal{R}_{\mu(s)\rightarrow\mu(t)}\otimes\mathcal{R}_{\mu(s)\rightarrow\mu(t)})(\cdot)(\mathcal{R}^{\top}_{\mu(s)\rightarrow\mu(t)}\otimes\mathcal{R}^{\top}_{\mu(s)\rightarrow\mu(t)})$, where $\otimes$ denotes the tensor product.
The proofs of examples 2.4-2.5 are given in Section 3 of the Supplementary Materials. For other manifolds, such as the hyperbolic manifold and the Grassmannian manifold [10], the parallel transformation $\boldsymbol{\Phi}$ can be derived similarly.
@@
where $\mu(\cdot)$ is the basis point function (BPF) on a Riemannian manifold, and $k_{t}(\cdot,\cdot)$ defined in (1)) is an intrinsic covariance function on a tangent space at $\mu(t)$. In this paper, we call (2) an intrinsic Gaussian distribution with the basepoint vector $\boldsymbol{\mu}_{n}$ and the covariance matrix $K$.
2.3 Intrinsic Gaussian process regression
Using the intrinsic Gaussian process, we develop an intrinsic Gaussian process regression model (iGPR) for the manifold-valued response $y\in\mathcal{M}$ and the vector-valued predictor $\boldsymbol{x}\in\Omega$ via an unknown regression function $f:\Omega\rightarrow\mathcal{M}$. Let $\Omega=\mathbb{R}^{Q}$ first. Later, we will extend $\Omega$ to other spaces. Given the training data $\mathcal{D}_{\mathcal{M}}:=\left\{\left(\boldsymbol{x}_{t_{i}},y_{t_{i}}\right)\mid\boldsymbol{x}_{t_{i}}\in\right.\left.\mathbb{R}^{Q},y_{t_{i}}\in\mathcal{M},i=1,\cdots,n,t_{i}\in\mathbb{T}\right\}$, where $\mathbb{T}$ is a compact set, the main idea of the model is to assume an intrinsic Gaussian process as the prior of $f(\cdot)$. We consider ${T}_{\mu(t^{*})}\mathcal{M}$ as the common tangent space. We will show later that the estimation results are invariant to the choice of this common tangent space. The joint distribution between the training outputs $\boldsymbol{y}=\left(y_{t_{1}},\cdots,y_{t_{n}}\right)$ and the test outputs $y_{t^{*}}$ at $\boldsymbol{x}_{t^{*}}$ is given by
@@ 2.3 Intrinsic Gaussian process regression
where $\mu^{*}=\mu\left(t^{*}\right)$, $\boldsymbol{\mu}=\left(\mu\left(t_{1}\right),\ldots,\mu\left(t_{{n}}\right)\right)$, $K_{t^{*}}(\boldsymbol{X},\boldsymbol{X})=\{k_{t^{*}}(\boldsymbol{x}_{t_{i}},\boldsymbol{x}_{t_{j}})\}_{1\leq i,j\leq n}$ and $\boldsymbol{X}=\left(\boldsymbol{x}_{t_{1}},\cdots,\boldsymbol{x}_{t_{n}}\right)$, $K_{t^{*}}\left(\boldsymbol{X},\boldsymbol{x}^{*}\right)=(k_{t^{*}}\left(\boldsymbol{x}_{t_{1}},\boldsymbol{x}^{*}\right),k_{t^{*}}\left(\boldsymbol{x}_{t_{2}},\boldsymbol{x}^{*}\right),\cdots,k_{t^{*}}\left(\boldsymbol{x}_{t_{n}},\boldsymbol{x}^{*}\right))$ and $k_{t^{*}}(\cdot,\cdot)$ is defined in (3), but with the common tangent space at $\mu(t^{*})$.
@@
In practice, we might need to work with a specific orthonormal bases for the tangent space. Therefore, we pick an orthogonal coordinate frame $\mathbf{E}$ for the common tangent space ${T}_{\mu(t^{*})}\mathcal{M}$, and for any $t\in\mathbb{T}$, $\mathbf{E}_{t}=\Gamma_{\mu(t^{*})\rightarrow\mu(t)}(\mathbf{E})$ is an orthogonal coordinate frame for the tangent space ${T}_{\mu(t)}\mathcal{M}$. Next, we will also show that the model estimation results are invariant to the choice of the orthogonal coordinate frame. Let $k_{\mathbf{E},t^{*}}$ be the representation of the intrinsic tangent space covariance function $k_{t^{*}}$ in the frame $\mathbf{E}$, and $\text{cov}_{\mathbf{E}}$ be the representation of the covariance function in frame $\mathbf{E}$. Thence, we have $k_{t^{*}}=\mathbf{E}^{\top}k_{\mathbf{E},t^{*}}\mathbf{E}$. And for any $t_{1},t_{2}\in\mathbf{T}$, we have $\text{cov}_{\mathbf{E}}(\Gamma_{\mu(t_{1})\rightarrow\mu(t^{*})}(\boldsymbol{\gamma}_{t_{1}}),\Gamma_{\mu(t_{2})\rightarrow\mu(t^{*})}(\boldsymbol{\gamma}_{t_{2}})):=k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}})$. Note that $k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}})$ is the matrix-valued function. To obtain a valid $k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}})$, a commonly used method is the intrinsic coregionalization model. This leads to $k_{\mathbf{E},t^{*}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}}):=k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t_{1}},\boldsymbol{x}_{t_{2}},\boldsymbol{\theta})\otimes\mathbf{B}\quad\text{with}\quad\mathbf{B}\in\mathbf{R}^{D\times D},$ where $k_{\mathbb{R}^{Q}}$ is a scalar kernel function, $\boldsymbol{\theta}$ is its kernel parameters, and $\mathbf{B}$ is a $D\times D$ coregionalization matrix that is positive semi-definite. Let matrix $K_{t^{*}}=(k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t_{i}},\boldsymbol{x}_{t_{j}},\boldsymbol{\theta}))_{n\times n}$, and $\boldsymbol{k}_{t^{*}}=(k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t^{*}},\boldsymbol{x}_{t_{1}},\boldsymbol{\theta}),k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t^{*}},\boldsymbol{x}_{t_{2}},\boldsymbol{\theta}),\cdots,k_{\mathbb{R}^{Q}}(\boldsymbol{x}_{t^{*}},\boldsymbol{x}_{t_{n}},\boldsymbol{\theta}))^{\top}$. Then predictive distribution of the test data $\boldsymbol{x}_{t^{*}}$ in tangent space of $\mu(t^{*})$ conditional on the training data $\mathcal{D}_{\mathcal{M}}$ is obtained as
@@
It follows that $\operatorname{Exp}(\mu(t^{*}),{\boldsymbol{\beta}}_{t^{*}}^{*})$ can be used as a point prediction of $y_{t^{*}}$. Although the predictive distribution $\mathcal{N}_{t^{*}}$ is not necessarily intrinsic Gaussian distribution, as $\boldsymbol{\beta}_{t^{*}}^{*}$ might be non-zero. As in [19], the distribution can be approximated using a intrinsic Gaussian distribution with the basepoint $\operatorname{Exp}_{{\mu}(t^{*})}\left(\boldsymbol{\beta}_{t^{*}}^{*}\right)$ and parallel transporting the tangent space covariance $\boldsymbol{\Sigma}_{t^{*}}^{*}$ to this basepoint along the corresponding geodesic. In addition, these hyperparameters $\boldsymbol{\theta}$ and $\mathbf{B}$ can be obtained by maximizing the logarithmic marginal likelihood. As there is no analytic solution, the recursive algorithm developed by [11, 8] can be used. A detailed computation procedure for iGPR is proved in Algorithm 1.
Algorithm 1 iGPR
@@
Next, we need to address the problem of how to choose the common tangent space. The process so far has been carried out in the tangent space of $\mu(t^{*})$. What would happen if we were to perform the analysis in the tangent space of other points along the BPF ? Theorem 2.6 below provides an answer: the posterior distribution $\mathcal{N}_{t}$ under the tangent space of other points $\mu(t)$ on the BPF, after applying a parallel transformation $\boldsymbol{\Phi}_{\mu(t)\rightarrow\mu(t^{*})}$ along the BPF, remains consistent with the posterior distribution $\mathcal{N}_{t^{*}}$ in the tangent space of $\mu(t^{*})$. The proof is given in Section 4 of the Supplementary Materials.
@@
Assuming that the BPF $\mu(t)$ is a smooth curve. For any $t\in\mathbf{T}$, let $\mathbf{A}$ be a frame for the tangent space of $\mu(t)$ and $\Gamma_{\mu(t)\rightarrow\mu(t^{*})}(\mathbf{A})=\mathbf{E}$. Then, we have ${\boldsymbol{\beta}}_{t^{*}}^{*}=\Gamma_{\mu(t)\rightarrow\mu(t^{*})}({\boldsymbol{\beta}}_{t}^{*});\quad\boldsymbol{\Sigma}_{t^{*}}^{*}=\boldsymbol{\Phi}_{\mu(t)\rightarrow\mu(t^{*})}(\boldsymbol{\Sigma}_{t}^{*}),$ where ${\boldsymbol{\beta}}_{t}^{*}$ and $\boldsymbol{\Sigma}_{t}^{*}$ represent the mean and variance of the predicted distribution $\mathcal{N}_{t}$ in the tangent space of $\mu(t)$.
Furthermore, the following Theorem 2.7 demonstrates that the predicted distribution is independent of the chosen frame. The proof is given in Section 5 of the Supplementary Materials.
@@
For any $t\in\mathbf{T}$, suppose $\mathbf{W}$ is another frame on the tangent spaces of $\mu(t)$. Change from $\mathbf{E}$ to $\mathbf{W}$ can be characterized by a unitary matrix $\mathbf{O}$, and $\mathbf{W}=\mathbf{O}^{\top}\mathbf{E}$. Then, we have ${\boldsymbol{\beta}}_{\mathbf{E},t}^{*}=\mathbf{O}{\boldsymbol{\beta}}_{\mathbf{W},t}^{*};\quad\overline{\boldsymbol{\Sigma}}^{*}_{\mathbf{E},t}=\mathbf{O}\overline{\boldsymbol{\Sigma}}^{*}_{\mathbf{W},t}\mathbf{O}^{\top}.$ This is, the mean and variance of the predicted distribution is independent of the frame.
The proposed iGPR uses the curvature information of the manifold by implementing parallel transport from different tangent spaces along the BPF $\mu(t)$ to a common tangent space, affectively merging the intrinsic geometric structure of the manifold. We have shown that the estimation for the iGPR model is unaffected by the selection of coordinate frames or common tangent space. In addition, we find an equivalent version of the proposed iGPR: parallel transporting the coordinate frame along the BPF results in a series of coordinate frames across different tangent spaces, leading to a representation of random elements in these frames within a Euclidean space $\mathbb{R}^{D}$. We also show that these representations retain consistency despite changes in the tangent spaces. Figure 2 gives a geometric illustration, where the black arrows represent coordinate frames obtained via parallel transport operator on different tangent spaces, and the orange solid lines stand for the representation of random elements after a parallel transport from the common tangent space to different tangent spaces. This results in the purple dashed lines connecting $y_{i}$, which is the curve predicted from the intrinsic Gaussian process regression model. The relative position between the orange solid line and coordinate axes in black remains unchanged, indicating that the choice of different common tangent space does not affect the model estimation. After representing the random elements of different tangent spaces using the aforementioned coordinate frame, they reside within the same Euclidean space and can be modeled by a vector-valued Gaussian process regression model.
Figure 2: Logarithm map of the manifold-valued data and schematic diagram of equivalent version of iGPR, where the blue solid line represents BPF.
We assumed a Euclidean predictor so far. The proposed method, however, can be extended to a broader framework that can model response variables residing on manifolds with mixed Euclidean or non-Euclidean predictors.
@@
By incorporating kernel functions specifically designed for non-Euclidean predictors within the iGPR model, our proposed method still works. Let the predictor $\boldsymbol{x}\in\Omega$, where $\Omega$ is a general space such as a metric space. As long as there is a kernel function $k_{\Omega}$ with symmetric and positive definite properties, the estimation and prediction computations are carried out similarly to those in Section 2.3 while $k_{\mathbb{R}^{Q}}$ is replaced with $k_{\Omega}$. For example, when $\Omega$ is a Riemannian manifold, one might select either the heat kernel [29, 30] or the Matérn kernel [7]. When $\Omega$ is a general metric space, the options include the graph kernel [15, 38] and the persistence weighted Gaussian kernel [21].
2.4 iGPR with observation noise
We didn’t consider noise in the model discussed in Section 2.3. In practical applications, data typically include random fluctuations and observational errors. By incorporating noise into the iGPR framework, the model can more accurately distinguish between signal and noise and then capture the underlying patterns. Furthermore, the inclusion of noise guarantees the positive definiteness of the covariance matrix, ensuring numerical stability. The iGPR model with error terms can be defined as